OpenStudy (anonymous):

Please help, I've been having trouble with these difficult problems and I'm stuck ... Solve for (x) 1. x+5=36 2. x+19=21 3. x+5+3=10 4. 19-x=12 5. 100-x+2=52 This math is just too advanced for me. Go through the process if you want!

OpenStudy (anonymous):

x + 5 = 36 subtract 5 from both sides to get x by itself x = 31 answer

OpenStudy (anonymous):

2advanced4u

OpenStudy (anonymous):

I will only do number 1, as an example for you to do the rest. $$x+5=36 \Rightarrow x = 36 -5 \Rightarrow x = 31$$ When I moved the 5 over the equal sign, it becomes the opposite value that it was. Hence \(-5\). @douglaswinslowcooper beat me to it. :-P

OpenStudy (anonymous):

I just don't get it ... sometimes I wonder if I'm cut out for Algebra 3. Why I even took the class is beyond me!

OpenStudy (anonymous):

If you live in America, this is actually pre-algebra, not Algebra "3". (I am assuming you pressed 3 instead of 2 on your keyboard.)

OpenStudy (anonymous):

That's racist.

OpenStudy (anonymous):

Yes I live in America. Got shipped here in a box.

OpenStudy (anonymous):

Uh, okay. Well I am willing to help you not give answer. So try to attempt the other questions, if you cannot then tag me and I will try to walk you through the process.

OpenStudy (anonymous):

Okay here are my attempts ... 2. x=30 3. x=4 4. x=9 5. x=61

OpenStudy (anonymous):

I think I got them right.

OpenStudy (anonymous):

Looks like I won't be needing help after all! Maybe people will think I'm worth something again! :) Thank you for giving me a chance!

OpenStudy (anonymous):

Hold on, you got them wrong. (At least number 2, I haven't checked the rest.) Explain how you got 30 for number 2?

OpenStudy (anonymous):

You Just do the opposite from each side For ex. X + 19=21 subtract 19 from each side X + 19=21 -19 -19 x = 2 so x=2

OpenStudy (anonymous):

I'm not understanding ... maybe it's the accent that you guys are typing with. Please understand that I am foreign. Allah hu ackbar!

OpenStudy (anonymous):

I think we are being pranked.

OpenStudy (anonymous):

oh no, i am sorry. i went to the bathroom and my older brother typed this :(

OpenStudy (anonymous):

please move on

OpenStudy (anonymous):

what a jerk

OpenStudy (anonymous):

I'm not a jerk, just creative.

OpenStudy (anonymous):